### v1

# Queryable Prototype Multiple Instance Learning with Vision-Language Models for Incremental Whole Slide Image Classification

Abstract
Whole Slide Image (WSI) classification has very significant applications in clinical pathology, e.g., tumor identification and cancer diagnosis. Currently, most research attention is focused on Multiple Instance Learning (MIL) using static WSI datasets. One of the most obvious weaknesses of these methods is that they cannot efficiently preserve and utilize previously learned knowledge. With any new data arriving, classification models are required to be re-trained on both previous and current new data. To overcome this shortcoming and break through traditional vision modality, this paper proposes the first Vision-Language-based framework with Queryable Prototype Multiple Instance Learning (QPMIL-VL) specially designed for incremental WSI classification. This framework mainly consists of two information processing branches: one is for generating bag-level features by prototype-guided aggregation of instance features, while the other is for enhancing class features through a combination of class ensemble, tunable vector and class similarity loss. The experiments on four public WSI datasets demonstrate that our QPMIL-VL framework is effective for incremental WSI classification and often significantly outperforms other compared methods, achieving state-of-the-art (SOTA) performance.
Introduction
Histopathology Whole Slide Image (WSI) is crucial for the diagnosis and treatment of tumor diseases (Bera et al. 2019; Song et al. 2023). To model gigapixel WSIs (e.g., $000\times 90,$ 000 × 90 , 000 pixels) for training clinical-grade models, many methods based on Multiple Instance Learning (MIL) have been extensively studied in the field of computational pathology (CPATH) (Campanella et al. 2019; Lu et al. 2021; Lin et al. 2023, 2024; Liu et al. 2024c). A shared characteristic of these methods is that they are specially designed for capturing a static data distribution (e.g., a given WSI dataset with two lung cancer subtypes), still following conventional machine learning paradigm. However, in the real-world, the distribution of WSI data could be dynamic due to the arrival of new datasets or the discovery of emerging cancer types (Van der Laak, Litjens, and Ciompi 2021; Derakhshani et al. 2022). As a result, a well-trained WSI classification model that is suitable for previous data often cannot adapt to new data. Therefore, the model must be re-trained on the entire dataset (including both previous and new data), leading to a significant increase in training costs. This presents the necessities of new methods well-tailored to dynamic data distributions.
Figure 1: (a) Existing visual modality framework for incremental WSI classification with buffer dependency; (b) our proposed Vision-Language-based framework with a queryable prototype pool and class feature enhancement.
Incremental Learning (IL) (Li and Hoiem 2017), also referred to as continual learning or lifelong learning, is exactly such approach to filling the aforementioned gap. As a new learning paradigm, it can encourage models to maintain the memory stability on old data distributions while adapting to a new one (Wang et al. 2024), thus mitigating the notorious catastrophic forgetting problem (McCloskey and Cohen 1989). Owing to such practical benefit, this new paradigm has attracted great attention and has shown strong potential for application in the real-world (Li et al. 2019; Caccia et al. 2021; Wang, Huang, and Hong 2022). Similarly, in CPATH, IL strategies are also noticed recently and show promising performance in learning from dynamic WSI data. Concretely, ConSlide (Huang et al. 2023a) is proposed for the first time to study incremental WSI classification, as illustrated in Fig. 1 (a). (1) Like most existing methods, ConSlide captures image-level features for WSI classification, relying solely on a traditional vision modality. (2) Moreover, to retain certain past data, an additional buffer is utilized to enable the model to review previous knowledge while learning new information. This may result in privacy concerns (Shokri and Shmatikov 2015) and higher computational costs, leaving a gap to the resource efficiency goal pursued in IL (Wang et al. 2024).
Recent works, e.g., PLIP (Huang et al. 2023b), Prov-GigaPath (Xu et al. 2024), and CONCH (Lu et al. 2024), have deeply demonstrated the significant potential of Vision-Language Models (VLMs) in CPATH. Breaking through traditional vision modality, this kind of new models could effectively exploit pathology features from both vision and language modalities. All of these motivate us to leverage foundational pathology VLMs to study incremental learning for WSI classification.
To extend the pure vision frameworks, we propose a new Vision-Language-based framework with Queryable Prototype Multiple Instance Learning (QPMIL-VL) for incremental WSI classification, as shown in Fig. 1 (b). Specifically, (1) in the vision branch, inspired by L2P (Wang et al. 2022), we design a Queryable Prototype MIL (QPMIL) module comprising a prototype pool and prototype-guided aggregation. This module encourages the model to incrementally learn a set of prototypes corresponding to each dataset through a prototype-query mechanism, effectively mitigating catastrophic forgetting without relying on extra buffer. (2) In the language branch, we propose a Class Feature Enhancement (CFE) module. We employ CFE to increase the diversity of text descriptions via a class ensemble approach and further refine it with tunable vector and a class similarity loss. Our experiments show that QPMIL-VL could often surpass other state-of-the-art (SOTA) methods by a large margin in incremental WSI classification tasks. The contributions of this work are as follows:
I) To our knowledge, beyond traditional pure vision frameworks, we propose the first Vision-Language-based framework for incremental WSI classification.
II) We devise a new Queryable Prototype Multiple Instance Learning (QPMIL) strategy to alleviate catastrophic forgetting. In this strategy, instance features are matched with a set of prototypes through querying to guide the generation of WSI bag-level feature.
III) Extensive comparison and ablation experiments are performed on four public WSI datasets. The results demonstrate the superiority of our QPMIL-VL over the existing methods.
Related Work
Multiple Instance Learning for WSI Classification
Multiple Instance Learning (MIL) focuses on learning from weakly-annotated data, where only an unknown subset of instances within each input bag is relevant to the label. Due to the unique data characteristics of WSI, the MIL paradigm is widely applied in WSI classification (Campanella et al. 2019; Li, Li, and Eliceiri 2021; Shao et al. 2021; Zhang et al. 2022; Zheng et al. 2023; Liu et al. 2024b). Among these methods, ABMIL (Ilse, Tomczak, and Welling 2018) proposes an attention-based instance feature aggregation. By leveraging ResNet for instance-level feature extraction, CLAM (Lu et al. 2021) introduces an interpretable weakly-supervised learning method, focused on data-efficient WSI processing. However, most of these studies focus on static MIL classification tasks and rarely consider the incremental WSI classifcation with MIL.
Incremental Learning
Recently, Incremental Learning (IL) has gained significant attention, aiming to enable deep models to continually acquire knowledge like humans. IL algorithms can be classified into three main categories. Regularization-based methods (Kirkpatrick et al. 2017; Li and Hoiem 2017) aim to reduce catastrophic forgetting by limiting changes to key parameters but often fall short of optimal results. Architecture-based methods (Rusu et al. 2016; Mallya and Lazebnik 2018; Li et al. 2019; Ke, Liu, and Huang 2020) train an independent module for each task but typically apply only to task-incremental learning scenario requiring task identity during inference. Additionally, although prevalent Rehearsal-based methods (Chaudhry et al. 2018, 2019; Prabhu, Torr, and Dokania 2020; Buzzega et al. 2020; Caccia et al. 2021) could achieve SOTA performance on various benchmarks (Parisi et al. 2019; Mai et al. 2022), these methods rely on additional buffer to store past data. When buffer size is limited, the performance of such these methods deteriorates (Cha, Lee, and Shin 2021). They will even become inapplicable when historical data is unavailable (Shokri and Shmatikov 2015). Currently, most of these studies focus more on natural images, not medical gigapixel WSIs, except ConSlide (Huang et al. 2023a).
Vision-Language Models
Recent research has made significant success in developing Vision-Language Models (VLMs). For example, CLIP (Radford et al. 2021) learns SOTA image representations, by the training on 400 million (image, text) pairs. Coca (Yu et al. 2022) employs contrastive and captioning losses to pre-train a foundation model comprising an image-text encoder-decoder. Additionally, some specialized VLMs like PLIP (Huang et al. 2023b), Prov-GigaPath (Xu et al. 2024) and CONCH (Lu et al. 2024) have been developed in CPATH. These VLMs need large datasets for pre-training, so they could often show excellent generalizability capabilities. Moreover, some VLMs combined with certain fine-tuning methods (Zhou et al. 2022; Yu et al. 2023; Gao et al. 2024) are widely transferred to various downstream tasks (Wang et al. 2023; Qu et al. 2024; Li et al. 2024; Liu et al. 2024a). At present, to our knowledge, the potential of VLMs for incremental WSI classification remains unexplored.
Preliminaries
Problem Formulation
Incremental WSI classification requires a model to continuously learn new knowledge from sequential tasks on non-stationary datasets, while retaining the knowledge learned from prior tasks. We define a sequence of WSI datasets $\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{T}\}$, in which the $t$-th dataset $\mathcal{D}_{t}=\{(\boldsymbol{x}_{i}^{t},y_{i}^{t})\}_{i=1}^{n_{t}}$ contains the tuples of sample $\boldsymbol{x}_{i}^{t}$ and its corresponding bag-level label $y_{i}^{t}$, where $1\leq t\leq T$ and $n_{t}$ is the sample number in $\mathcal{D}_{t}$. When training is made on current dataset $\mathcal{D}_{t_{c}}$, the data from previous datasets (i.e., $\mathcal{D}_{1},\cdots,\mathcal{D}_{t_{c}-1}$) may be limited or even unavailable.
For a given WSI sample $\boldsymbol{x}_{i}^{t}$, in class-incremental learning scenario, the model only uses $\boldsymbol{x}_{i}^{t}$ to predict its bag-level label $y_{i}^{t}$. In contrast, in task-incremental learning scenario, the model combines sample $\boldsymbol{x}_{i}^{t}$ with its corresponding task identity $t$ together to predict label $y_{i}^{t}$.
VLM Pre-training and Inference
In our work, we choose a SOTA VLM in pathology, i.e., CONCH (Lu et al. 2024), as our image and text encoders.
Pre-training. CONCH uses the contrastive learning on diverse histopathology images, biomedical text, and over $1.17$ million image-caption pairs (not including TCGA Datasets). Based on CoCa (Yu et al. 2022) framework, CONCH combines an image encoder $\boldsymbol{E}_{\text{img}}$ with $f(\cdot,\theta)$, a text encoder $\boldsymbol{E}_{\text{txt}}$ with $g(\cdot,\phi)$, and a multi-modal fusion decoder.
Zero-shot Patch Inference. Since CONCH is pre-trained on patch-level WSI image-text pairs, after training is completed, zero-shot inference can be fulfilled on patch-level samples. Specifically, given a patch $\boldsymbol{p}$ and class text token set $\{\boldsymbol{t}_{c}\}_{c=1}^{C}$, where $C$ denotes total class number. The classification can be conducted based on the cosine similarity between patch feature $f(\boldsymbol{p},\theta)$ and class feature $g(\boldsymbol{t}_{c},\phi)$.
Figure 2: The framework of QPMIL-VL. The prompts in the prototype pool enable an efficient incremental learning process by gradually capturing the visual feature descriptions of instance prototypes present in the sequential WSI datasets.
Proposed Method
To effectively fulfil incremental WSI classification, on traditional MIL framework, we introduce Vision-Language Model and propose a queryable prototype MIL framework, as shown in Fig. 2. It consists of two branches, one for generating bag-level feature (QPMIL), and the other for enhancing class text feature (CFE). In inference, the features from both branches will be used to compute classification probability.
Queryable Prototype Multiple Instance Learning (QPMIL)
As usual, one of the most important objectives for WSI classification is to generate discriminative bag-level feature representation. However, unlike traditional static classification, our incremental classification task requires that all subsequent model learning specifically avoids causing catastrophic impacts on the tasks learned earlier.
Moreover, it has been proved that the instances in different WSIs, even in different WSI datasets, could often appear high similarities, in terms of cell shape, staining response, and tissue arrangement (Song et al. 2024).Therefore, similar instances, even if they are from different datasets, could be clustered into the same instance prototype. In view of this, to generate more effective WSI bag-level feature, we introduce VLM and instance-level feature prototypes, designing a queryable prototype pool and prototype-guided aggregation.
Prototype Pool. Specifically, the prototype pool consists of $M$ (key, prompt) prototype pairs $\{(\boldsymbol{k}_{i},\boldsymbol{P}_{i})\}_{i=1}^{M},\boldsymbol{k}_{i}\in\mathbb{R}^{D_{f}},\boldsymbol{P}_{i}\in\mathbb{R}^{L_{\boldsymbol{P}}\times D_{e}}$, where $D_{f}$ is the output feature dimension of encoder $\boldsymbol{E}_{\text{img}}$ and $\boldsymbol{E}_{\text{txt}}$, $D_{e}$ represents the embedding dimension of $\boldsymbol{E}_{\text{txt}}$ and $L_{\boldsymbol{P}}$ indicates the total number of learnable vectors in $\boldsymbol{P}_{i}$. In a prototype pair, the key works as a prototype identifier, used by WSIs for querying, while the prompt acts as a prototype descriptor, utilized to describe the specific visual feature of instance prototype within WSIs.
Given a WSI sample $\boldsymbol{x}$ from $\mathcal{D}_{t}$, it is first divided into $n$ non-overlapping patches $\{\boldsymbol{p}_{i}\}_{i=1}^{n}$, i.e., instances. Then, all instances are encoded by image encoder $\boldsymbol{E}_{\text{img}}$ with $f(\cdot,\theta)$, generating $n$ instance features $\boldsymbol{Z}\in\mathbb{R}^{n\times D_{f}}$, as shown in Fig. 2. Next, we apply MaxPooling to features $\boldsymbol{Z}$ to obtain the query vector of sample $\boldsymbol{x}$, i.e., $\boldsymbol{z}=MaxPooling(\boldsymbol{Z})\in\mathbb{R}^{D_{f}}$. MeanPooling is also tried and its experimental result is provided in our Supplementary Material.
Subsequently, a query metric function $q:\mathbb{R}^{D_{f}}\times\mathbb{R}^{D_{f}}\to\mathbb{R}$ is defined to evaluate how close the query vector $\boldsymbol{z}$ matches the key of each prototype pair. We employs the cosine distance for this evaluation. Finally, we are able to find the most matched prototype keys through
$\mathcal{K}_{\boldsymbol{x}}=Top\text{-}N^{min}\{q(\boldsymbol{z},\boldsymbol{k}_{i})\}_{i=1}^{M},$ (1)
where $Top\text{-}N^{min}$ denotes the operation of choosing the top-$N$ prototype keys with minimal $\boldsymbol{z}$-key distances. However, in prototype matching on different WSI datasets by Eq. 1, we observe that almost all samples tend to choose the same set of $N$ prototypes from pool. This could usually lead to catastrophic forgetting in incremental learning. To effectively avoid this tendency, we define a penalty factor $p_{i}^{t_{c}}$ to optimize prototype choosing:
$\mathcal{K}_{\boldsymbol{x}}&=Top\text{-}N^{min}\{q(\boldsymbol{z},\boldsymbol{k}_{i})\cdot p_{i}^{t_{c}}\}_{i=1}^{M}\\ P_{t_{c}}&=\frac{1}{t_{c}-1}\sum_{t=1}^{t_{c}-1}F_{t},t_{c}\geqslant 2,\end{aligned}\right.$ (2)
where $t_{c}$ is the index of current WSI dataset $\mathcal{D}_{t_{c}}$, $F_{t}=[f_{1}^{t},\cdots,f_{M}^{t}]$ denotes the matching frequency table of $M$ prototype keys on dataset $\mathcal{D}_{t}$ and $p_{i}^{t_{c}}$ represents the $i$-th element of penalty table $P_{t_{c}}$. In essence, this penalty factor amplifies the distances of $\boldsymbol{z}$ to all the prototype keys that have high frequency matching on those datasets prior to $\mathcal{D}_{t_{c}}$.
Prototype-guided Aggregation. After key matching, the prompts $\mathcal{P}_{\boldsymbol{x}}=\{\boldsymbol{P}_{i}^{\boldsymbol{x}}\}_{i=1}^{N}$ of top-$N$ prototype pairs will be further fed into text encoder $\boldsymbol{E}_{\text{txt}}$ to generate $N$ prototype features $\boldsymbol{F}_{p}\in\mathbb{R}^{N\times D_{f}}$.
Next, we compute the cosine similarity matrix $\boldsymbol{S}\in\mathbb{R}^{n\times N}$ between instance features $\boldsymbol{Z}$ and prototype features $\boldsymbol{F}_{p}$. Since each column in $\boldsymbol{S}$ exactly corresponds to a prototype, a Softmax normalization along column direction is performed to obtain a weighted aggregation matrix $\boldsymbol{W}\in\mathbb{R}^{n\times N}$. Finally, we obtain the prototype-guided bag-level features through $\boldsymbol{W}$, and their average yields the bag-level feature $\boldsymbol{f}_{b}\in\mathbb{R}^{D_{f}}$ for the WSI sample $\boldsymbol{x}$. The procedure can be represented as
$\boldsymbol{f}_{b}=mean(Softmax(\boldsymbol{Z}\times{\boldsymbol{F}_{p}}^{\top})^{\top}\times\boldsymbol{Z}).$ (3)
Ideally, in incremental training, different datasets will match a non-overlapped set of prototype pairs. In practice, some prototype pairs could often be shared with low frequencies by different datasets. Our method doesn’t impose a strict independence of matched prototype pairs. Instead, it allows WSI datasets to choose their optimal prototypes, providing more flexibility and adaptability.
Class Feature Enhancement (CFE)
As shown in Figure 2, class text feature enhancing branch is designed to generate an enhanced class feature $\boldsymbol{f^{\prime}}_{\text{txt}}$.
Class Ensemble. For a specific subtype of cancer, such as “lung adenocarcinoma”, it can also be expressed as “adenocarcinoma of the lung”. Therefore, we adopt class ensemble to obtain better class feature. For each cancer subtype, we combine several common class names and templates to obtain $m$ different class text descriptions. All text descriptions are encoded by text encoder $\boldsymbol{E}_{\text{txt}}$ with $g(\cdot,\phi)$ to obtain initial class feature matrix $\boldsymbol{F}_{\text{txt}}\in\mathbb{R}^{m\times D_{f}}$, and by averaging $\boldsymbol{F}_{\text{txt}}$, we get the class feature $\boldsymbol{f}_{\text{txt}}\in\mathbb{R}^{D_{f}}$ for this subtype. The entire calculation process can be expressed as
$\boldsymbol{f}_{\text{txt}}=mean(g(\boldsymbol{T},\phi)),$ (4)
where $\boldsymbol{T}\in\mathbb{R}^{m\times L}$ represents the text tokens corresponding to the text descriptions and $L$ is the length of text token.
Tunable Vector. To further refine class feature $\boldsymbol{f}_{\text{txt}}$, referring to the TaskRes (Yu et al. 2023), we design a tunable (i.e., learnable) vector $\boldsymbol{v}\in\mathbb{R}^{D_{f}}$ of the same dimension for each class. The final enhanced class feature $\boldsymbol{f^{\prime}}_{\text{txt}}\in\mathbb{R}^{D_{f}}$ is calculated through
$\boldsymbol{f^{\prime}}_{\text{txt}}=\boldsymbol{f}_{\text{txt}}+\alpha\cdot\boldsymbol{v},$ (5)
where $\alpha$ (default constant $0.5$) is an additional amplitude factor to control the superposition scale of vector $\boldsymbol{v}$ on feature $\boldsymbol{f}_{\text{txt}}$.
Prediction and Optimization
Class Prediction. The probability of predicting the WSI sample $\boldsymbol{x}$ as the class $y$ can be computed as
$p(y|\boldsymbol{x})=\frac{e^{\tau\cdot\langle\boldsymbol{f}_{b},[\boldsymbol{f^{\prime}}_{\text{txt}}]_{y}\rangle}}{\sum_{c=1}^{C}e^{\tau\cdot\langle\boldsymbol{f}_{b},[\boldsymbol{f^{\prime}}_{\text{txt}}]_{c}\rangle}},$ (6)
[…]
Optimization Objective. To appropriately reduce the similarity between features of different classes, we firstly define a class similarity loss, which is calculated through
$\mathcal{L}_{S}=\frac{2}{C\cdot(C-1)}\sum_{i=1}^{C}\sum_{j=i+1}^{C}\langle[\boldsymbol{f^{\prime}}_{\text{txt}}]_{i},[\boldsymbol{f^{\prime}}_{\text{txt}}]_{j}\rangle+1.$ (7)
Moreover, there are another two losses. One is classification loss. Based on Eq. 6, it is computed by
$\mathcal{L}_{C}=-\text{log}\frac{e^{\tau\cdot\langle\boldsymbol{f}_{b},[\boldsymbol{f^{\prime}}_{\text{txt}}]_{y}\rangle}}{\sum_{c=1}^{C}e^{\tau\cdot\langle\boldsymbol{f}_{b},[\boldsymbol{f^{\prime}}_{\text{txt}}]_{c}\rangle}}.$ (8)
IL Type Method ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$) Params (M) Masked ACC ($\uparrow$)
Baseline JointTrain (Upper) 0.908 $\pm$ 0.022 1.000 $\pm$ $0$ 0.502 0.937 $\pm$ 0.022
FineTune (Lower) 0.308 $\pm$ 0.045 0.336 $\pm$ 0.045 0.841 $\pm$ 0.054 -0.841 $\pm$ 0.054 0.844 $\pm$ 0.038
Regularization-based EWC (Kirkpatrick et al. 2017) 0.309 $\pm$ 0.051 0.338 $\pm$ 0.055 0.836 $\pm$ 0.064 -0.836 $\pm$ 0.064 0.502 0.840 $\pm$ 0.035
LwF (Li and Hoiem 2017) 0.378 $\pm$ 0.081 0.412 $\pm$ 0.083 0.731 $\pm$ 0.116 -0.731 $\pm$ 0.116 0.916 $\pm$ 0.031
Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.436 $\pm$ 0.058 0.477 $\pm$ 0.070 0.670 $\pm$ 0.072 -0.670 $\pm$ 0.072 0.502 0.879 $\pm$ 0.045
ER-ACE/30 (Caccia et al. 2021) 0.666 $\pm$ 0.049 0.732 $\pm$ 0.051 0.075 $\pm$ 0.057 -0.070 $\pm$ 0.059 0.917 $\pm$ 0.023
DER++/30 (Buzzega et al. 2020) 0.749 $\pm$ 0.055 0.823 $\pm$ 0.059 0.219 $\pm$ 0.059 -0.214 $\pm$ 0.062 0.904 $\pm$ 0.029
ER/30 (Chaudhry et al. 2019) 0.790 $\pm$ 0.040 0.870 $\pm$ 0.049 0.186 $\pm$ 0.044 -0.186 $\pm$ 0.044 0.894 $\pm$ 0.036
ConSlide/30 (Huang et al. 2023a) 0.659 $\pm$ 0.022 - 0.076 $\pm$ 0.030 -0.075 $\pm$ 0.030 39.446 0.861 $\pm$ 0.017
ER/100 (Chaudhry et al. 2019) 0.827 $\pm$ 0.028 0.910 $\pm$ 0.034 0.143 $\pm$ 0.034 -0.142 $\pm$ 0.032 0.502 0.918 $\pm$ 0.028
Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.616 $\pm$ 0.056 0.677 $\pm$ 0.071 0.285 $\pm$ 0.059 -0.285 $\pm$ 0.059 0.111 0.844 $\pm$ 0.026
MI-Zero (Lu et al. 2023) 0.839 $\pm$ 0.034 0.927 $\pm$ 0.044 - - - 0.909 $\pm$ 0.019
PQMIL-VL (ours) 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014 0.365 0.930 $\pm$ 0.018
Table 1: Main results in class-incremental learning scenario on forward-order training. The best performances are highlighted as bold. “Params” represents the number of learnable parameters. Gray numbers are directly cited from published paper. “/30” and “/100” mean buffer $30$ WSIs and buffer $100$ WSIs, respectively. We present the Masked ACC reflecting task-incremental learning scenario for reference.
The other is matching loss to make the matched keys $\mathcal{K}_{\boldsymbol{x}}=\{\boldsymbol{k}_{i}^{\boldsymbol{x}}\}_{i=1}^{N}$ closer to the query vector $\boldsymbol{z}$ of WSI sample $\boldsymbol{x}$. It is defined as
$\mathcal{L}_{M}=\frac{1}{N}\sum_{i=1}^{N}q(\boldsymbol{z},\boldsymbol{k}_{i}^{\boldsymbol{x}}).$ (9)
To optimize our QPMIL-VL model in training, based on the three losses above, total loss $\mathcal{L}_{O}$ is computed by
$\mathcal{L}_{O}=\mathcal{L}_{C}+\lambda\cdot\mathcal{L}_{M}+\beta\cdot\mathcal{L}_{S},$ (10)
where balance factors $\lambda$ and $\beta$ are set to constant $0.5$. Therefore, our optimization objective is to minimize $\mathcal{L}_{O}$.
Experiments
Following the setup of previous representative work (Van de Ven and Tolias 2019), we mainly evaluate QPMIL-VL in the class-incremental learning scenario and additionally report the results in the task-incremental scenario.
Experimental Details
Datasets. Following ConSlide (Huang et al. 2023a), we use four public WSI datasets from The Cancer Genome Atla (TCGA) repository: non-small cell lung carcinoma (NSCLC), invasive breast carcinoma (BRCA), renal cell carcinoma (RCC) and esophageal carcinoma (ESCA). With the arrival of each dataset, we add two additional classes to the model for training and evaluation.
We use CLAM (Lu et al. 2021) to crop non-overlapping $256\times 256$ patches from the segmented tissue at 10×10$\times$10 × magnification. Then, pre-trained image encoder $\boldsymbol{E}_{\text{img}}$ in CONCH (Lu et al. 2024) is used to extract instance features.
Evaluation Metrics. We assess all incremental learning methods with Average Accuracy (ACC), Upper-bound Ratio, Forgetting and Backward Transfer (BWT) (Lopez-Paz and Ranzato 2017; Hayes et al. 2018; Fini et al. 2022). Besides, we provide the Masked Average Accuracy (Masked ACC) metric, which is used for reference only to evaluate performance in task-incremental learning scenario by masking the logits of irrelevant classes. All results are reported by the mean and standard deviation of ten-fold cross-validation.
[…]
Refer to our Supplementary Material for more details.
Comparison Results
Compared Methods. It is often necessary to compare with the upper and lower bounds of classification methods. We employ the classic ABMIL (Ilse, Tomczak, and Welling 2018) as the aggregator for the vision branch, serving as a Baseline. Its performances by jointly training on all datasets (JointTrain, non-incremental) are taken as upper bound, and those by naively fine-tuning sequential datasets (FineTune) are taken as lower bound. In addition, we compare Regularization-based methods: EWC (Kirkpatrick et al. 2017) and LwF (Li and Hoiem 2017); Rehearsal-based methods: A-GEM (Chaudhry et al. 2018), ER (Chaudhry et al. 2019), DER++ (Buzzega et al. 2020), ER-ACE (Caccia et al. 2021) and ConSlide (Huang et al. 2023a); and Pre-trained VLM-based methods: AttriCLIP (Wang et al. 2023) and MI-Zero (Lu et al. 2023). To ensure a fair comparison, all compared methods are adapted to VLM. Specially, they use the same instance features and text encoder $\boldsymbol{E}_{\text{txt}}$ in CONCH.
IL Type Method ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$) Masked ACC ($\uparrow$)
Baseline JointTrain (Upper) 0.908 $\pm$ 0.022 1.000 $\pm$ $0$ 0.937 $\pm$ 0.022
FineTune (Lower) 0.234 $\pm$ 0.008 0.262 $\pm$ 0.010 0.927 $\pm$ 0.023 -0.927 $\pm$ 0.023 0.803 $\pm$ 0.060
Regularization-based EWC (Kirkpatrick et al. 2017) 0.235 $\pm$ 0.011 0.264 $\pm$ 0.011 0.928 $\pm$ 0.020 -0.928 $\pm$ 0.020 0.833 $\pm$ 0.069
LwF (Li and Hoiem 2017) 0.236 $\pm$ 0.016 0.265 $\pm$ 0.012 0.908 $\pm$ 0.030 -0.908 $\pm$ 0.030 0.900 $\pm$ 0.041
Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.536 $\pm$ 0.047 0.591 $\pm$ 0.053 0.527 $\pm$ 0.067 -0.527 $\pm$ 0.067 0.872 $\pm$ 0.026
ER-ACE/30 (Caccia et al. 2021) 0.703 $\pm$ 0.049 0.777 $\pm$ 0.053 0.281 $\pm$ 0.062 -0.279 $\pm$ 0.063 0.889 $\pm$ 0.041
DER++/30 (Buzzega et al. 2020) 0.684 $\pm$ 0.055 0.755 $\pm$ 0.063 0.310 $\pm$ 0.069 -0.307 $\pm$ 0.072 0.910 $\pm$ 0.043
ER/30 (Chaudhry et al. 2019) 0.644 $\pm$ 0.028 0.711 $\pm$ 0.033 0.387 $\pm$ 0.050 -0.386 $\pm$ 0.049 0.901 $\pm$ 0.035
ConSlide/30 (Huang et al. 2023a) 0.499 $\pm$ 0.025 - 0.058 $\pm$ 0.032 -0.021 $\pm$ 0.039 0.854 $\pm$ 0.039
Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.694 $\pm$ 0.058 0.766 $\pm$ 0.061 0.207 $\pm$ 0.063 -0.207 $\pm$ 0.063 0.861 $\pm$ 0.019
MI-Zero (Lu et al. 2023) 0.839 $\pm$ 0.034 0.927 $\pm$ 0.044 - - 0.909 $\pm$ 0.019
PQMIL-VL (ours) 0.859 $\pm$ 0.032 0.946 $\pm$ 0.028 0.064 $\pm$ 0.031 -0.064 $\pm$ 0.031 0.925 $\pm$ 0.018
Table 2: Main results in class-incremental learning scenario on reverse-order training.
[…]
(1) Our QPMIL-VL method is consistently superior to all other methods, with parameters only slightly larger than the smallest one. PQMIL-VL achieves an ACC of $0.890$$\pm$ $0.021$ with an Upper-bound Ratio as high as $0.982$$\pm$ $0.032$, winning SOTA performance metrics. Its Upper-bound Ratio is even $5.5\%$ higher than that of MI-Zero, the second-best method. Additionally, our model has a small learnable parameter size of $0.365$M, which is only slightly larger than the smallest size of $0.111$M in AttriCLIP. However, in terms of performance metrics, our method is far superior to the lowest-parameter method, i.e., AttriCLIP. Besides, in task-incremental experiments, our Masked ACC could reach up to $0.930$$\pm$ $0.018$, which is also significantly higher than any achieved by other compared methods.
(2) Two regularization-based methods are almost ineffective in WSI class-incremental learning scenario. For example, as a regularization-based method, LwF only achieves an ACC of $0.378$$\pm$ $0.081$, which is slightly higher than the lower bound of $0.308$$\pm$ $0.045$, resulting in a very low Upper-bound Ratio of $0.412$$\pm$ $0.083$. Meawhile, EWC also shows almost ineffective in incremental classification.
(3) Most rehearsal-based and pre-trained VLM-based methods could achieve good classification results. The rehearsal-based DER++ acquires an ACC of $0.749$$\pm$ $0.055$ with an Upper-bound Ratio of $0.823$$\pm$ $0.059$, and the pre-trained VLM-based MI-Zero obtains an ACC of $0.839$$\pm$ $0.034$ with an Upper-bound Ratio of $0.927$$\pm$ $0.044$. Their Upper-bound Ratio results exceed $80\%$. Therefore, these two types of methods are believed to be more effective than the former, consistent with natural image observations. Notably, although AttriCLIP performs well in the natural image-based IL setting, its performance in this task is lower than some rehearsal-based methods, reflecting the unique challenges of analyzing gigapixel WSI. We consider the main reasons are the poor bag-level feature (obtained through MaxPooling) and the overly similar class features derived from the same set of prompts.
QPMIL CFE ACC ($\uparrow$) Forgetting ($\downarrow$)
Tunable Vector Class Ensemble Class Similarity Loss
0.308 $\pm$ 0.045 0.841 $\pm$ 0.054
$\checkmark$ 0.720 $\pm$ 0.029 0.158 $\pm$ 0.025
$\checkmark$ $\checkmark$ 0.747 $\pm$ 0.039 0.149 $\pm$ 0.037
$\checkmark$ $\checkmark$ $\checkmark$ 0.874 $\pm$ 0.033 0.052 $\pm$ 0.030
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.890 $\pm$ 0.021 0.027 $\pm$ 0.014
Table 3: Ablation studies on main components.
Ablation ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$)
w/o key 0.675 $\pm$ 0.029 0.742 $\pm$ 0.045 0.332 $\pm$ 0.034 -0.332 $\pm$ 0.034
w/o matching penalty 0.659 $\pm$ 0.029 0.724 $\pm$ 0.037 0.356 $\pm$ 0.048 -0.356 $\pm$ 0.048
QPMIL-VL 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014
Table 4: Ablation studies on prototype pool.
[…]
(1) Our QPMIL-VL method is consistently superior to all other methods in both class-incremental and task-incremental learning scenarios, maintaining a SOTA position. (2) Regularization-based methods are still ineffective in class-incremental learning scenario due to suffering from severe catastrophic forgetting. (3) Most rehearsal-based and pre-trained VLM-based methods often achieve good classification performance. Except for MI-Zero and our PQMIL-VL, the best ACC among these two types of methods reaches up to $0.703$$\pm$ $0.049$ with an Upper-bound Ratio of $0.777$$\pm$ $0.053$, achieved by ER-ACE.
Comparison between Forward-order and Reverse-order Training. Additionally, we observe that on reverse-order training, the performance of most methods is noticeably lower than on forward-order training. For example, our PQMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$. In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Table 1). One possible explanation to this decline is that in incremental learning, later datasets can affect earlier ones. ESCA, with only $150$ slides, is the first to be learned, while NSCLC, with $965$ slides, is the last. NSCLC is dominant in the total number of training samples, which could cause a greater influence on the earlier datasets, leading to more forgetting.
Ablation and Analysis
Ablation Studies. The following are several ablation analyses of QPMIL-VL.
I) on Main Components To analyze the impact of the main components on performance, we conduct a set of ablation studies on forward-order training. The results are shown in Table 3, where the first row represents our baseline (FineTune). By comparison, two obvious findings can be observed as follows.
(1) Each component in QPMIL-VL contributes to the performance improvement. For example, starting from the baseline, QPMIL can significantly boost the ACC from $0.308$$\pm$ $0.045$ to $0.720$$\pm$ $0.029$. Similarly, incrementally applying Class Similarity Loss raises the ACC from $0.874$$\pm$ $0.033$ to $0.890$$\pm$ $0.021$. (2) Among all components, both QPMIL and Class Ensemble have a more significant impact on performance than the other two. This is because the prototype query mechanism in QPMIL effectively mitigates catastrophic forgetting, while the Class Ensemble significantly optimizes class feature encoding in the VLM by increasing the diversity of class text descriptions. These comparisons imply that each component is effective.
II) on Prototype Pool Furthermore, we conduct another set of ablation studies specifically on the queryable prototype pool. Our experimental results are shown in Table 4.
(1)w/o key” indicates the removal of the key from prototype pair, implying that incremental training always uses the same set of prompts. In this situation, the knowledge from different datasets is encoded into the same set of prompts, leading to catastrophic forgetting and a significant performance decline (ACC dropping to $0.675$$\pm$ $0.029$). (2)w/o matching penalty” presents that only Eq. 1 is used for matching prototype pairs. Under this ablation, the ACC metric decreases to $0.659$$\pm$ $0.029$ because different datasets tend to match the same set of prototypes during training, which consequently results in catastrophic forgetting. Only when both key and matching penalty are applied together in our QPMIL-VL does the ACC reach its highest value of $0.890$$\pm$ $0.021$. This ablation demonstrates that both are beneficial for incremental classification.
Further Performance Analysis. Additionally, we conduct additional experiments to further analyze our method.
I) Effect of Hyper-parameters As shown in Fig. 3, there are three important parameters in the prototype pool. The results indicate that inappropriate prototype pool capacity ((a): $N=7$, (b): $M<20$) may result in knowledge interference and too few prompts ((a): $N=1$) makes it difficult to describe the various instance prototype visual features present in WSIs.
Figure 3: (a) ACC w.r.t length of each prompt ($L_{\boldsymbol{P}}$) and size of matching keys ($N$), size of prototype pool $M=20$; (b) ACC w.r.t $M$, $L_{\boldsymbol{P}}=24$ and $N=5$.
Figure 4: (a) Prototype key matching frequency histogram; (b) prototype feature visualization.
II) Prototype Key Matching To analyze the matching results of keys on incremental datasets, we compute the matching frequency of each prototype pair in the first fold of experiments, as illustrated in Fig. 4 (a). It can be observed that each dataset tends to stably match $5$ keys. Moreover, the $5$ keys predominantly matched by different datasets are often non-overlapping due to the matching penalty.
III) Prototype Feature Visualization As depicted in Fig. 4 (b), we present the distribution visualization of $200$ prototype features using t-SNE (Van der Maaten and Hinton 2008). These features are derived from the ten-fold cross-validation experiments conducted on four datasets, i.e., $200=10\times 4\times 5$. Their distributions show that most prototypes learned from the same dataset exhibit clear clustering properties. From another perspective, they also imply that our model can learn distinct prototype features from different datasets. Additionally, from Fig. 4 (a) and (b), we observe a small number of outliers. They can be regarded as shared prototypes.
IV) Class Feature Enhancement Analysis To show the effectiveness of CFE, we visualize four classes of WSI samples and their corresponding class features, as shown in Fig. 5. More experimental results can be found in our Supplementary Material.
From the alteration in feature similarity, we can observe that the class features are noticeably improved by the enhancement. For example, in the case of the class sample ESCA-ESCC, the initial cosine similarity between the center point feature and the class feature is only $0.078$, but it increases to $0.447$ after enhancement, resulting in an improvement of $0.369$. This alteration suggests that the CFE is effective in enhancing the quality of class features.
Figure 5: Class features visualization, only one class per dataset. Red $\star$ is the center, i.e., the average of features, for the instances of interest (orange points, determined by the cosine similarity between instance features and learned prototype features). Green $\circ$ and $\bullet$ denote the cosine similarities before and after feature enhancement, respectively.
Conclusions
This paper proposes the first Vision-Language-based framework for incremental WSI classification, QPMIL-VL, breaking through traditional vision modality. By a pool of learnable prototype pairs, it aggregates the instance features from pre-trained image encoder to generate bag-level feature under the guidance of matched prototype prompts. Class probability is predicted by the Consine similarity between bag-level feature and enhanced class feature. Extensive experiments demonstrate that our method and its components are effective. It obviously outperforms other methods in incremental WSI classification, achieving SOTA results.
Appendix A Supplementary Material
A. Dataset Details
The benchmark consists of four public WSI datasets from TCGA repository: non-small cell lung carcinoma (NSCLC), invasive breast carcinoma (BRCA), renal cell carcinoma (RCC) and esophageal carcinoma (ESCA), as shown in Table 5.
Dataset Tumor Type Cases Slides
NSCLC Lung adenocarcinoma (LUAD) 449 512
Lung squamous cell carcinoma (LUSC) 419 453
BRCA Invasive ductal (IDC) 713 761
Invasive lobular carcinoma (ILC) 178 191
RCC Clear cell renal cell carcinoma (CCRCC) 493 499
Papillary renal cell carcinoma (PRCC) 245 269
ESCA Esophageal adenocarcinoma (ESAD) 64 64
Esophageal squamous cell carcinoma (ESCC) 84 86
Table 5: The statistics of incremental WSI analysis benchmark.
B. Metric Details
After completing the incremental training, we can obtain an Accuracy performance matrix, as illustrated in Table 6. Subsequently, the relevant metrics reported in our experiments are calculated using this performance matrix.
After Training Test on
Dataset $1$ Dataset $2$cdots Dataset $T$
Dataset $1$ $R_{1,1}$cdots -
Dataset $2$ $R_{2,1}$ $R_{2,2}$cdots -
$\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$
Dataset $T$ $R_{T,1}$ $R_{T,2}$cdots $R_{T,T}$
JointTrain (Upper) $R_{joint,1}$ $R_{joint,2}$cdots $R_{joint,T}$
Table 6: The Accuracy performance matrix during incremental training.
ACC. Average Accuracy (ACC) represents the model’s average performance across all datasets, which is calculated by
$ACC=\frac{1}{T}\sum_{t=1}^{T}R_{T,t}.$ (11)
Upper-bound Ratio. Upper-bound Ratio enables an incrementally trained model to be compared relative to a jointly trained model (upper bound). It can be computed by
$Upper\text{-}bound\ Ratio=\frac{1}{T}\sum_{t=1}^{T}\frac{R_{T,t}}{R_{joint,t}}.$ (12)
Forgetting. Forgetting quantifies how much knowledge the model has forgotten about previous datasets, which is defined as
$Forgetting=\frac{1}{T-1}\sum_{t=1}^{T-1}\max_{i\in\{1,\cdots,T\}}R_{i,t}-R_{T,t}.$ (13)
BWT. Backward Transfer (BWT) reflects the model’s ability to mitigate catastrophic forgetting (i.e., memory stability). The calculation formula is as follows:
$BWT=\frac{1}{T-1}\sum_{t=1}^{T-1}R_{T,t}-R_{t,t}.$ (14)
C. More Details of Class Ensemble
Table 7 provides all the class names and templates used in our implementation of the class ensemble. They are adapted from those used in CONCH.
Template Dataset Tumor Type ClassName
“ClassName” “a photomicrograph showing ClassName” “a photomicrograph of ClassName” “an image of ClassName” “an image showing ClassName” “an example of ClassName” “ClassName is shown” “this is ClassName” “there is ClassName” “a histopathological image showing ClassName” “a histopathological image of ClassName” “a histopathological photograph of ClassName” “a histopathological photograph showing ClassName” “shows ClassName” “presence of ClassName” “ClassName is present” “an H&E stained image of ClassName” “an H&E stained image showing ClassName” “an H&E image showing ClassName” “an H&E image of ClassName” “ClassName, H&E stain” “ClassName, H&E” NSCLC LUAD “lung adenocarcinoma”
“adenocarcinoma of the lung”
“LUAD”
LUSC “lung squamous cell carcinoma”
“squamous cell carcinoma of the lung”
“LUSC”
BRCA IDC “invasive ductal carcinoma”
“breast invasive ductal carcinoma”
“invasive ductal carcinoma of the breast”
“invasive carcinoma of the breast, ductal pattern”
“breast IDC”
ILC “invasive lobular carcinoma”
“breast invasive lobular carcinoma”
“invasive lobular carcinoma of the breast”
“invasive carcinoma of the breast, lobular pattern”
“breast ILC”
RCC CCRCC “clear cell renal cell carcinoma”
“renal cell carcinoma, clear cell type”
“renal cell carcinoma of the clear cell type”
“clear cell RCC”
PRCC “papillary renal cell carcinoma”
“renal cell carcinoma, papillary type”
“renal cell carcinoma of the papillary type”
“papillary RCC”
ESCA ESAD “esophageal adenocarcinoma”
“adenocarcinoma of the esophageal”
“ESAD”
ESCC “esophageal squamous cell carcinoma”
“squamous cell carcinoma of the esophageal”
“ESCC”
Table 7: The class names and templates used in our class ensemble.
D. Analysis on Accuracy Performance Matrix
The performance variations of deep models across different datasets during incremental learning cannot be thoroughly analyzed solely by relying on final metrics such as ACC. Therefore, we present detailed evaluation results of QPMIL-VL on both forward-order and reverse-order incremental training in Tables 8 and 9, respectively. Through our analysis, there are three findings.
After Training Test on
NSCLC BRCA RCC ESCA
NSCLC 0.906 $\pm$ 0.027 - - -
BRCA 0.895 $\pm$ 0.027 0.878 $\pm$ 0.024 - -
RCC 0.892 $\pm$ 0.024 0.876 $\pm$ 0.025 0.931 $\pm$ 0.022 -
ESCA 0.837 $\pm$ 0.046 0.868 $\pm$ 0.029 0.929 $\pm$ 0.023 0.928 $\pm$ 0.083
JointTrain (Upper) 0.891 $\pm$ 0.041 0.908 $\pm$ 0.029 0.921 $\pm$ 0.045 0.913 $\pm$ 0.055
Table 8: The Accuracy performance matrix of QPMIL-VL on forward-order training.
After Training Test on
ESCA RCC BRCA NSCLC
ESCA 0.967 $\pm$ 0.045 - - -
RCC 0.967 $\pm$ 0.045 0.932 $\pm$ 0.029 - -
BRCA 0.955 $\pm$ 0.060 0.930 $\pm$ 0.030 0.873 $\pm$ 0.030 -
NSCLC 0.807 $\pm$ 0.092 0.913 $\pm$ 0.039 0.860 $\pm$ 0.028 0.856 $\pm$ 0.029
JointTrain (Upper) 0.913 $\pm$ 0.055 0.921 $\pm$ 0.045 0.908 $\pm$ 0.029 0.891 $\pm$ 0.041
Table 9: The Accuracy performance matrix of QPMIL-VL on reverse-order training.
(1) Forgetting phenomena are commonly observed in the old datasets. On forward-order training, the Accuracy of NSCLC decreases from $0.906$$\pm$ $0.027$ to $0.837$$\pm$ $0.046$, while on reverse-order training, the Accuracy of ESCA drops from $0.967$$\pm$ $0.045$ to $0.807$$\pm$ $0.092$. This reveals the prevalent phenomenon of forgetting during the incremental learning process.
(2) Our QPMIL-VL demonstrates competitiveness in static WSI classification when compared to the baseline (JointTrain). For example, on forward-order training, after QPMIL-VL is trained on the fourth dataset, ESCA, it reaches an Accuracy of $0.928$$\pm$ $0.083$, which is higher than the $0.913$$\pm$ $0.055$ of JointTrain. However, due to forgetting caused by incremental learning, QPMIL-VL shows a noticeable drop in performance on earlier datasets, so its overall performance ends up being lower than that of JointTrain.
(3) The primary reason for the performance degradation on reverse-order training compared to forward-order training lies in the greater impact of learning the fourth dataset on the performance of the first dataset. After training on the fourth dataset, the Accuracy of the first dataset decreases by $0.055$ ($0.892$$\pm$ $0.024$to $0.837$$\pm$ $0.046$) and $0.148$ ($0.955$$\pm$ $0.060$to $0.807$$\pm$ $0.092$) on forward-order and reverse-order training, respectively. During reverse training, NSCLS significantly affects the performance of ESCA, this may be because the NSCLS sample size ($965$) is about six times larger than that of ESCA ($150$), making it easier for NSCLS to interfere with ESCA. In contrast, during forward training, ESCA has a much smaller impact on NSCLS.
E. Query Vector Generation Strategies
MaxPooling vs MeanPooling. To investigate the impact of different pooling operations on the performance of QPMIL-VL when generating query vector, we conduct comparative experiments using MaxPooling and MeanPooling. Table 10 presents the experimental results.
The results show that QPMIL-VL performs better overall with MaxPooling compared to MeanPooling, particularly on forward-order training (the ACC increases from $0.877$$\pm$ $0.034$ to $0.890$$\pm$ $0.021$). These results reflect that the query vectors obtained through MaxPooling could possess greater discriminative ability across different datasets (refer to Fig. 6).
Training Order Pooling Type ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$) Masked ACC ($\uparrow$)
MeanPooling 0.877 $\pm$ 0.034 0.966 $\pm$ 0.039 0.045 $\pm$ 0.019 -0.044 $\pm$ 0.019 0.929 $\pm$ 0.027
Forward-order MaxPooling 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014 0.930 $\pm$ 0.018
MeanPooling 0.859 $\pm$ 0.042 0.946 $\pm$ 0.042 0.070 $\pm$ 0.054 -0.070 $\pm$ 0.055 0.935 $\pm$ 0.015
Reverse-order MaxPooling 0.859 $\pm$ 0.032 0.946 $\pm$ 0.028 0.064 $\pm$ 0.031 -0.064 $\pm$ 0.031 0.925 $\pm$ 0.018
Table 10: The comparative results of MaxPooling and MeanPooling on both forward-order and reverse-order training.
F. Prototype Visualization on Whole Slide Images
The pre-defined prompts in the prototype pool are designed to describe the visual features of the instance prototypes of interest present in WSIs. To validate this design principle, we calculate the cosine similarity between the prompt features (i.e., prototype features) learned by QPMIL-VL and the instance features, and use this similarity as the basis to determine the prototype category to which each instance belongs. The experimental results are presented in Fig. 789 and 10.
Figure 6: The visualization of the query vectors from the four datasets obtained through MeanPooling (left) and MaxPooling (right).
By observing the results, we can clearly see that prompts can correctly identify the target instance prototypes (i.e., the lesion areas in the WSIs). Additionally, different prompts can locate different instance prototypes to some extent. This experimental phenomenon is consistent with our expectations.
Figure 7: Prototype visualization on Whole Slide Images (NSCLC).
Figure 8: Prototype visualization on Whole Slide Images (BRCA).
Figure 9: Prototype visualization on Whole Slide Images (RCC).
Figure 10: Prototype visualization on Whole Slide Images (ESCA).
G. More Analysis on Class Feature Enhancement
Since in QPMIL-VL, class features guide the learning of WSI bag-level features, we further validate the effectiveness of the CFE module by visualizing the distribution of the bag-level features learned by the model before and after using the CFE module in four different incremental learning stages. The results are shown in Fig. 11.
From these results, we could find that our CFE makes i) intra-class bag-level features more compact (dashed box) and ii) inter-class ones more distinguishable (solid box). This finding suggests that the class features used for supervising bag-level feature learning are improved.
Figure 11: Bag-level features visualization with or without our Class Feature Enhancement (CFE) in four incremental learning stages.

### v2

@@
 The experiments on four public WSI datasets demonstrate that our QPMIL-VL framework is effective for incremental WSI classification and often significantly outperforms other compared methods, achieving state-of-the-art (SOTA) performance.
+Our source code is publicly available at https://github.com/can-can-ya/QPMIL-VL.
 Introduction
-Histopathology Whole Slide Image (WSI) is crucial for the diagnosis and treatment of tumor diseases (Bera et al.
-2019; Song et al.
+Histopathology Whole Slide Image (WSI) is crucial for the diagnosis and treatment of tumor diseases (Song et al.
 2023).
@@
 2021; Lin et al.
-2023, 2024; Liu et al.
+2023; Liu et al.
 2024c).
@@
 As a result, a well-trained WSI classification model that is suitable for previous data often cannot adapt to new data.
-Therefore, the model must be re-trained on the entire dataset (including both previous and new data), leading to a significant increase in training costs.
+Therefore, it must be re-trained on the entire dataset (including both previous and new data), leading to a significant increase in training costs.
 This presents the necessities of new methods well-tailored to dynamic data distributions.
@@
 2023b), Prov-GigaPath (Xu et al.
-2024), and CONCH (Lu et al.
+2024) and CONCH (Lu et al.
 2024), have deeply demonstrated the significant potential of Vision-Language Models (VLMs) in CPATH.
@@
 In this strategy, instance features are matched with a set of prototypes through querying to guide the generation of WSI bag-level feature.
-III) Extensive comparison and ablation experiments are performed on four public WSI datasets. The results demonstrate the superiority of our QPMIL-VL over the existing methods.
+III) Extensive experiments are performed on four public WSI datasets. The results demonstrate the superiority of our QPMIL-VL over the existing methods.
 Related Work
@@
 Prototype Pool.
-Specifically, the prototype pool consists of $M$ (key, prompt) prototype pairs $\{(\boldsymbol{k}_{i},\boldsymbol{P}_{i})\}_{i=1}^{M},\boldsymbol{k}_{i}\in\mathbb{R}^{D_{f}},\boldsymbol{P}_{i}\in\mathbb{R}^{L_{\boldsymbol{P}}\times D_{e}}$, where $D_{f}$ is the output feature dimension of encoder $\boldsymbol{E}_{\text{img}}$ and $\boldsymbol{E}_{\text{txt}}$, $D_{e}$ represents the embedding dimension of $\boldsymbol{E}_{\text{txt}}$ and $L_{\boldsymbol{P}}$ indicates the total number of learnable vectors in $\boldsymbol{P}_{i}$.
+Specifically, the prototype pool consists of $M$ (key, prompt) prototype pairs $\{(\boldsymbol{k}_{i},\boldsymbol{P}_{i})\}_{i=1}^{M}$ initialized using a Normal Distribution, $\boldsymbol{k}_{i}\in\mathbb{R}^{D_{f}},\boldsymbol{P}_{i}\in\mathbb{R}^{L_{\boldsymbol{P}}\times D_{e}}$, where $D_{f}$ is the output feature dimension of encoder $\boldsymbol{E}_{\text{img}}$ and $\boldsymbol{E}_{\text{txt}}$, $D_{e}$ represents the embedding dimension of $\boldsymbol{E}_{\text{txt}}$ and $L_{\boldsymbol{P}}$ indicates the total number of learnable vectors in $\boldsymbol{P}_{i}$.
 In a prototype pair, the key works as a prototype identifier, used by WSIs for querying, while the prompt acts as a prototype descriptor, utilized to describe the specific visual feature of instance prototype within WSIs.
@@
 where $Top\text{-}N^{min}$ denotes the operation of choosing the top-$N$ prototype keys with minimal $\boldsymbol{z}$-key distances.
-However, in prototype matching on different WSI datasets by Eq.
+However, in prototype matching on different WSI datasets during incremental training by Eq.
 1, we observe that almost all samples tend to choose the same set of $N$ prototypes from pool.
@@
 Baseline JointTrain (Upper) 0.908 $\pm$ 0.022 1.000 $\pm$ $0$ 0.502 0.937 $\pm$ 0.022
-FineTune (Lower) 0.308 $\pm$ 0.045 0.336 $\pm$ 0.045 0.841 $\pm$ 0.054 -0.841 $\pm$ 0.054 0.844 $\pm$ 0.038
-Regularization-based EWC (Kirkpatrick et al. 2017) 0.309 $\pm$ 0.051 0.338 $\pm$ 0.055 0.836 $\pm$ 0.064 -0.836 $\pm$ 0.064 0.502 0.840 $\pm$ 0.035
-LwF (Li and Hoiem 2017) 0.378 $\pm$ 0.081 0.412 $\pm$ 0.083 0.731 $\pm$ 0.116 -0.731 $\pm$ 0.116 0.916 $\pm$ 0.031
-Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.436 $\pm$ 0.058 0.477 $\pm$ 0.070 0.670 $\pm$ 0.072 -0.670 $\pm$ 0.072 0.502 0.879 $\pm$ 0.045
-ER-ACE/30 (Caccia et al. 2021) 0.666 $\pm$ 0.049 0.732 $\pm$ 0.051 0.075 $\pm$ 0.057 -0.070 $\pm$ 0.059 0.917 $\pm$ 0.023
-DER++/30 (Buzzega et al. 2020) 0.749 $\pm$ 0.055 0.823 $\pm$ 0.059 0.219 $\pm$ 0.059 -0.214 $\pm$ 0.062 0.904 $\pm$ 0.029
-ER/30 (Chaudhry et al. 2019) 0.790 $\pm$ 0.040 0.870 $\pm$ 0.049 0.186 $\pm$ 0.044 -0.186 $\pm$ 0.044 0.894 $\pm$ 0.036
+FineTune (Lower) 0.308 $\pm$ 0.045 ∗∗∗ 0.336 $\pm$ 0.045 0.841 $\pm$ 0.054 -0.841 $\pm$ 0.054 0.844 $\pm$ 0.038 ∗∗∗
+Regularization-based EWC (Kirkpatrick et al. 2017) 0.309 $\pm$ 0.051 ∗∗∗ 0.338 $\pm$ 0.055 0.836 $\pm$ 0.064 -0.836 $\pm$ 0.064 0.502 0.840 $\pm$ 0.035 ∗∗∗
+LwF (Li and Hoiem 2017) 0.378 $\pm$ 0.081 ∗∗∗ 0.412 $\pm$ 0.083 0.731 $\pm$ 0.116 -0.731 $\pm$ 0.116 0.916 $\pm$ 0.031
+Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.436 $\pm$ 0.058 ∗∗∗ 0.477 $\pm$ 0.070 0.670 $\pm$ 0.072 -0.670 $\pm$ 0.072 0.502 0.879 $\pm$ 0.045 ∗∗
+ER-ACE/30 (Caccia et al. 2021) 0.666 $\pm$ 0.049 ∗∗∗ 0.732 $\pm$ 0.051 0.075 $\pm$ 0.057 -0.070 $\pm$ 0.059 0.917 $\pm$ 0.023
+DER++/30 (Buzzega et al. 2020) 0.749 $\pm$ 0.055 ∗∗∗ 0.823 $\pm$ 0.059 0.219 $\pm$ 0.059 -0.214 $\pm$ 0.062 0.904 $\pm$ 0.029 ∗∗
+ER/30 (Chaudhry et al. 2019) 0.790 $\pm$ 0.040 ∗∗∗ 0.870 $\pm$ 0.049 0.186 $\pm$ 0.044 -0.186 $\pm$ 0.044 0.894 $\pm$ 0.036 ∗∗
 ConSlide/30 (Huang et al. 2023a) 0.659 $\pm$ 0.022 - 0.076 $\pm$ 0.030 -0.075 $\pm$ 0.030 39.446 0.861 $\pm$ 0.017
-ER/100 (Chaudhry et al. 2019) 0.827 $\pm$ 0.028 0.910 $\pm$ 0.034 0.143 $\pm$ 0.034 -0.142 $\pm$ 0.032 0.502 0.918 $\pm$ 0.028
-Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.616 $\pm$ 0.056 0.677 $\pm$ 0.071 0.285 $\pm$ 0.059 -0.285 $\pm$ 0.059 0.111 0.844 $\pm$ 0.026
-MI-Zero (Lu et al. 2023) 0.839 $\pm$ 0.034 0.927 $\pm$ 0.044 - - - 0.909 $\pm$ 0.019
-PQMIL-VL (ours) 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014 0.365 0.930 $\pm$ 0.018
+ER/100 (Chaudhry et al. 2019) 0.827 $\pm$ 0.028 ∗∗∗ 0.910 $\pm$ 0.034 0.143 $\pm$ 0.034 -0.142 $\pm$ 0.032 0.502 0.918 $\pm$ 0.028
+Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.616 $\pm$ 0.056 ∗∗∗ 0.677 $\pm$ 0.071 0.285 $\pm$ 0.059 -0.285 $\pm$ 0.059 0.111 0.844 $\pm$ 0.026 ∗∗∗
+MI-Zero (Lu et al. 2023) 0.839 $\pm$ 0.034 ∗∗∗ 0.927 $\pm$ 0.044 - - - 0.909 $\pm$ 0.019 ∗∗
+QPMIL-VL (ours) 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014 0.365 0.930 $\pm$ 0.018
 Table 1: Main results in class-incremental learning scenario on forward-order training.
@@
 We present the Masked ACC reflecting task-incremental learning scenario for reference.
+/∗∗/∗∗∗ denote there are significant differences (paired t-test p-value $<$ $0.05$/$0.01$/$0.001$) between the best performance and others.
 The other is matching loss to make the matched keys $\mathcal{K}_{\boldsymbol{x}}=\{\boldsymbol{k}_{i}^{\boldsymbol{x}}\}_{i=1}^{N}$ closer to the query vector $\boldsymbol{z}$ of WSI sample $\boldsymbol{x}$.
@@
 $\mathcal{L}_{M}=\frac{1}{N}\sum_{i=1}^{N}q(\boldsymbol{z},\boldsymbol{k}_{i}^{\boldsymbol{x}}).$ (9)
-To optimize our QPMIL-VL model in training, based on the three losses above, total loss $\mathcal{L}_{O}$ is computed by
-$\mathcal{L}_{O}=\mathcal{L}_{C}+\lambda\cdot\mathcal{L}_{M}+\beta\cdot\mathcal{L}_{S},$ (10)
-where balance factors $\lambda$ and $\beta$ are set to constant $0.5$. Therefore, our optimization objective is to minimize $\mathcal{L}_{O}$.
+To optimize our QPMIL-VL model in training, based on the three losses above, total loss $\mathcal{L}_{T}$ is computed by
+$\mathcal{L}_{T}=\mathcal{L}_{C}+\lambda\cdot\mathcal{L}_{M}+\beta\cdot\mathcal{L}_{S},$ (10)
+where balance factors $\lambda$ and $\beta$ are set to constant $0.5$. Therefore, our optimization objective is to minimize $\mathcal{L}_{T}$.
 Experiments
-Following the setup of previous representative work (Van de Ven and Tolias 2019), we mainly evaluate QPMIL-VL in the class-incremental learning scenario and additionally report the results in the task-incremental scenario.
+Following the setup of previous representative work (Van de Ven and Tolias 2019), we mainly evaluate QPMIL-VL in the class-incremental learning scenario and additionally report the results in the task-incremental learning scenario.
 Experimental Details
@@
 Compared Methods.
-It is often necessary to compare with the upper and lower bounds of classification methods.
+It is necessary to compare with the upper and lower bounds of IL methods.
 We employ the classic ABMIL (Ilse, Tomczak, and Welling 2018) as the aggregator for the vision branch, serving as a Baseline.
@@
 2023).
-To ensure a fair comparison, all compared methods are adapted to VLM.
+To ensure a fair comparison, all methods are adapted to VLM.
 Specially, they use the same instance features and text encoder $\boldsymbol{E}_{\text{txt}}$ in CONCH.
@@
 Baseline JointTrain (Upper) 0.908 $\pm$ 0.022 1.000 $\pm$ $0$ 0.937 $\pm$ 0.022
-FineTune (Lower) 0.234 $\pm$ 0.008 0.262 $\pm$ 0.010 0.927 $\pm$ 0.023 -0.927 $\pm$ 0.023 0.803 $\pm$ 0.060
-Regularization-based EWC (Kirkpatrick et al. 2017) 0.235 $\pm$ 0.011 0.264 $\pm$ 0.011 0.928 $\pm$ 0.020 -0.928 $\pm$ 0.020 0.833 $\pm$ 0.069
-LwF (Li and Hoiem 2017) 0.236 $\pm$ 0.016 0.265 $\pm$ 0.012 0.908 $\pm$ 0.030 -0.908 $\pm$ 0.030 0.900 $\pm$ 0.041
-Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.536 $\pm$ 0.047 0.591 $\pm$ 0.053 0.527 $\pm$ 0.067 -0.527 $\pm$ 0.067 0.872 $\pm$ 0.026
-ER-ACE/30 (Caccia et al. 2021) 0.703 $\pm$ 0.049 0.777 $\pm$ 0.053 0.281 $\pm$ 0.062 -0.279 $\pm$ 0.063 0.889 $\pm$ 0.041
-DER++/30 (Buzzega et al. 2020) 0.684 $\pm$ 0.055 0.755 $\pm$ 0.063 0.310 $\pm$ 0.069 -0.307 $\pm$ 0.072 0.910 $\pm$ 0.043
-ER/30 (Chaudhry et al. 2019) 0.644 $\pm$ 0.028 0.711 $\pm$ 0.033 0.387 $\pm$ 0.050 -0.386 $\pm$ 0.049 0.901 $\pm$ 0.035
+FineTune (Lower) 0.234 $\pm$ 0.008 ∗∗∗ 0.262 $\pm$ 0.010 0.927 $\pm$ 0.023 -0.927 $\pm$ 0.023 0.803 $\pm$ 0.060 ∗∗∗
+Regularization-based EWC (Kirkpatrick et al. 2017) 0.235 $\pm$ 0.011 ∗∗∗ 0.264 $\pm$ 0.011 0.928 $\pm$ 0.020 -0.928 $\pm$ 0.020 0.833 $\pm$ 0.069 ∗∗∗
+LwF (Li and Hoiem 2017) 0.236 $\pm$ 0.016 ∗∗∗ 0.265 $\pm$ 0.012 0.908 $\pm$ 0.030 -0.908 $\pm$ 0.030 0.900 $\pm$ 0.041
+Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.536 $\pm$ 0.047 ∗∗∗ 0.591 $\pm$ 0.053 0.527 $\pm$ 0.067 -0.527 $\pm$ 0.067 0.872 $\pm$ 0.026 ∗∗∗
+ER-ACE/30 (Caccia et al. 2021) 0.703 $\pm$ 0.049 ∗∗∗ 0.777 $\pm$ 0.053 0.281 $\pm$ 0.062 -0.279 $\pm$ 0.063 0.889 $\pm$ 0.041
+DER++/30 (Buzzega et al. 2020) 0.684 $\pm$ 0.055 ∗∗∗ 0.755 $\pm$ 0.063 0.310 $\pm$ 0.069 -0.307 $\pm$ 0.072 0.910 $\pm$ 0.043
+ER/30 (Chaudhry et al. 2019) 0.644 $\pm$ 0.028 ∗∗∗ 0.711 $\pm$ 0.033 0.387 $\pm$ 0.050 -0.386 $\pm$ 0.049 0.901 $\pm$ 0.035
 ConSlide/30 (Huang et al. 2023a) 0.499 $\pm$ 0.025 - 0.058 $\pm$ 0.032 -0.021 $\pm$ 0.039 0.854 $\pm$ 0.039
-Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.694 $\pm$ 0.058 0.766 $\pm$ 0.061 0.207 $\pm$ 0.063 -0.207 $\pm$ 0.063 0.861 $\pm$ 0.019
+Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.694 $\pm$ 0.058 ∗∗∗ 0.766 $\pm$ 0.061 0.207 $\pm$ 0.063 -0.207 $\pm$ 0.063 0.861 $\pm$ 0.019 ∗∗∗
 MI-Zero (Lu et al. 2023) 0.839 $\pm$ 0.034 0.927 $\pm$ 0.044 - - 0.909 $\pm$ 0.019
-PQMIL-VL (ours) 0.859 $\pm$ 0.032 0.946 $\pm$ 0.028 0.064 $\pm$ 0.031 -0.064 $\pm$ 0.031 0.925 $\pm$ 0.018
+QPMIL-VL (ours) 0.859 $\pm$ 0.032 0.946 $\pm$ 0.028 0.064 $\pm$ 0.031 -0.064 $\pm$ 0.031 0.925 $\pm$ 0.018
 Table 2: Main results in class-incremental learning scenario on reverse-order training.
@@
 (1) Our QPMIL-VL method is consistently superior to all other methods, with parameters only slightly larger than the smallest one.
-PQMIL-VL achieves an ACC of $0.890$$\pm$ $0.021$ with an Upper-bound Ratio as high as $0.982$$\pm$ $0.032$, winning SOTA performance metrics.
+QPMIL-VL achieves an ACC of $0.890$$\pm$ $0.021$ with an Upper-bound Ratio as high as $0.982$$\pm$ $0.032$, winning SOTA performance metrics.
 Its Upper-bound Ratio is even $5.5\%$ higher than that of MI-Zero, the second-best method.
@@
 (3) Most rehearsal-based and pre-trained VLM-based methods often achieve good classification performance.
-Except for MI-Zero and our PQMIL-VL, the best ACC among these two types of methods reaches up to $0.703$$\pm$ $0.049$ with an Upper-bound Ratio of $0.777$$\pm$ $0.053$, achieved by ER-ACE.
+Except for MI-Zero and our QPMIL-VL, the best ACC among these two types of methods reaches up to $0.703$$\pm$ $0.049$ with an Upper-bound Ratio of $0.777$$\pm$ $0.053$, achieved by ER-ACE.
 Comparison between Forward-order and Reverse-order Training.
 Additionally, we observe that on reverse-order training, the performance of most methods is noticeably lower than on forward-order training.
-For example, our PQMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$.
+For example, our QPMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$.
 In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Table 1).
@@
 3, there are three important parameters in the prototype pool.
-The results indicate that inappropriate prototype pool capacity ((a): $N=7$, (b): $M<20$) may result in knowledge interference and too few prompts ((a): $N=1$) makes it difficult to describe the various instance prototype visual features present in WSIs.
+The results indicate that inappropriate prototype pool capacity ((a): $N=7$, (b): $M<20$) may result in knowledge interference and too few prompts ((a): $N=1$) make it difficult to describe the various instance prototype visual features present in WSIs.
 Figure 3: (a) ACC w.r.t length of each prompt ($L_{\boldsymbol{P}}$) and size of matching keys ($N$), size of prototype pool $M=20$; (b) ACC w.r.t $M$, $L_{\boldsymbol{P}}=24$ and $N=5$.
@@
 Red $\star$ is the center, i.e., the average of features, for the instances of interest (orange points, determined by the cosine similarity between instance features and learned prototype features).
+Blue points are non-interest instance features.
 Green $\circ$ and $\bullet$ denote the cosine similarities before and after feature enhancement, respectively.
@@
 By a pool of learnable prototype pairs, it aggregates the instance features from pre-trained image encoder to generate bag-level feature under the guidance of matched prototype prompts.
-Class probability is predicted by the Consine similarity between bag-level feature and enhanced class feature.
+Class probability is predicted by the cosine similarity between bag-level feature and enhanced class feature.
 Extensive experiments demonstrate that our method and its components are effective.
 It obviously outperforms other methods in incremental WSI classification, achieving SOTA results.
+Acknowledgments
+This work is supported by the National Natural Science Foundation of China (NSFC) under Grant No.62476049.
 Appendix A Supplementary Material

### v3

@@
 Class Feature Enhancement (CFE)
-As shown in Figure 2, class text feature enhancing branch is designed to generate an enhanced class feature $\boldsymbol{f^{\prime}}_{\text{txt}}$.
+As shown in Fig. 2, class text feature enhancing branch is designed to generate an enhanced class feature $\boldsymbol{f^{\prime}}_{\text{txt}}$.
 Class Ensemble.
@@
 $\mathcal{L}_{C}=-\text{log}\frac{e^{\tau\cdot\langle\boldsymbol{f}_{b},[\boldsymbol{f^{\prime}}_{\text{txt}}]_{y}\rangle}}{\sum_{c=1}^{C}e^{\tau\cdot\langle\boldsymbol{f}_{b},[\boldsymbol{f^{\prime}}_{\text{txt}}]_{c}\rangle}}.$ (8)
-IL Type Method ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$) Params (M) Masked ACC ($\uparrow$)
+IL Type Method ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$) Params Masked ACC ($\uparrow$)
 Baseline JointTrain (Upper) 0.908 $\pm$ 0.022 1.000 $\pm$ $0$ 0.502 0.937 $\pm$ 0.022
@@
 The best performances are highlighted as bold.
-“Params” represents the number of learnable parameters.
+“Params” represents the number of learnable parameters (M).
 Gray numbers are directly cited from published paper.
@@
 (2) Two regularization-based methods are almost ineffective in WSI class-incremental learning scenario.
-For example, as a regularization-based method, LwF only achieves an ACC of $0.378$$\pm$ $0.081$, which is slightly higher than the lower bound of $0.308$$\pm$ $0.045$, resulting in a very low Upper-bound Ratio of $0.412$$\pm$ $0.083$.
-Meawhile, EWC also shows almost ineffective in incremental classification.
+For example, as a regularization-based method, LwF only achieves an ACC of $0.378$$\pm$ $0.081$, which is slightly higher than the lower bound of $0.308$$\pm$ $0.045$.
 (3) Most rehearsal-based and pre-trained VLM-based methods could achieve good classification results.
@@
 QPMIL CFE ACC ($\uparrow$) Forgetting ($\downarrow$)
-Tunable Vector Class Ensemble Class Similarity Loss
+TV CE CSL
 0.308 $\pm$ 0.045 0.841 $\pm$ 0.054
@@
 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.890 $\pm$ 0.021 0.027 $\pm$ 0.014
-Table 3: Ablation studies on main components.
-Ablation ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$)
-w/o key 0.675 $\pm$ 0.029 0.742 $\pm$ 0.045 0.332 $\pm$ 0.034 -0.332 $\pm$ 0.034
-w/o matching penalty 0.659 $\pm$ 0.029 0.724 $\pm$ 0.037 0.356 $\pm$ 0.048 -0.356 $\pm$ 0.048
-QPMIL-VL 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014
+Table 3: Ablation studies on main components (TV: Tunable Vector, CE: Class Ensemble, CSL: Class Similarity Loss).
+Ablation ACC ($\uparrow$) Forgetting ($\downarrow$)
+w/o key 0.675 $\pm$ 0.029 0.332 $\pm$ 0.034
+w/o matching penalty 0.659 $\pm$ 0.029 0.356 $\pm$ 0.048
+QPMIL-VL 0.890 $\pm$ 0.021 0.027 $\pm$ 0.014
 Table 4: Ablation studies on prototype pool.
@@
 (3) Most rehearsal-based and pre-trained VLM-based methods often achieve good classification performance.
-Except for MI-Zero and our QPMIL-VL, the best ACC among these two types of methods reaches up to $0.703$$\pm$ $0.049$ with an Upper-bound Ratio of $0.777$$\pm$ $0.053$, achieved by ER-ACE.
 Comparison between Forward-order and Reverse-order Training.
@@
 For example, our QPMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$.
-In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Table 1).
-One possible explanation to this decline is that in incremental learning, later datasets can affect earlier ones.
+In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Tab.
+1).
+One possible explanation to this decline is that in IL, later datasets can affect earlier ones.
 ESCA, with only $150$ slides, is the first to be learned, while NSCLC, with $965$ slides, is the last.
@@
 I) on Main Components To analyze the impact of the main components on performance, we conduct a set of ablation studies on forward-order training.
-The results are shown in Table 3, where the first row represents our baseline (FineTune).
-By comparison, two obvious findings can be observed as follows.
+The results are shown in Tab.
+3, where the first row represents our baseline (FineTune).
+Two obvious findings can be observed as follows.
 (1) Each component in QPMIL-VL contributes to the performance improvement.
@@
 These comparisons imply that each component is effective.
-II) on Prototype Pool Furthermore, we conduct another set of ablation studies specifically on the queryable prototype pool. Our experimental results are shown in Table 4.
+II) on Prototype Pool Furthermore, we conduct another set of ablation studies specifically on the queryable prototype pool. Our experimental results are shown in Tab. 4.
 (1)w/o key” indicates the removal of the key from prototype pair, implying that incremental training always uses the same set of prompts.
@@
 A. Dataset Details
-The benchmark consists of four public WSI datasets from TCGA repository: non-small cell lung carcinoma (NSCLC), invasive breast carcinoma (BRCA), renal cell carcinoma (RCC) and esophageal carcinoma (ESCA), as shown in Table 5.
+The benchmark consists of four public WSI datasets from TCGA repository: non-small cell lung carcinoma (NSCLC), invasive breast carcinoma (BRCA), renal cell carcinoma (RCC) and esophageal carcinoma (ESCA), as shown in Tab.
+5.
 Dataset Tumor Type Cases Slides
@@
 B. Metric Details
-After completing the incremental training, we can obtain an Accuracy performance matrix, as illustrated in Table 6.
+After completing the incremental training, we can obtain an Accuracy performance matrix, as illustrated in Tab.
+6.
 Subsequently, the relevant metrics reported in our experiments are calculated using this performance matrix.
@@
 C. More Details of Class Ensemble
-Table 7 provides all the class names and templates used in our implementation of the class ensemble. They are adapted from those used in CONCH.
+Tab. 7 provides all the class names and templates used in our implementation of the class ensemble. They are adapted from those used in CONCH.
 Template Dataset Tumor Type ClassName
@@
 ESCA 0.837 $\pm$ 0.046 0.868 $\pm$ 0.029 0.929 $\pm$ 0.023 0.928 $\pm$ 0.083
-JointTrain (Upper) 0.891 $\pm$ 0.041 0.908 $\pm$ 0.029 0.921 $\pm$ 0.045 0.913 $\pm$ 0.055
+JointTrain 0.891 $\pm$ 0.041 0.908 $\pm$ 0.029 0.921 $\pm$ 0.045 0.913 $\pm$ 0.055
 Table 8: The Accuracy performance matrix of QPMIL-VL on forward-order training.
@@
 NSCLC 0.807 $\pm$ 0.092 0.913 $\pm$ 0.039 0.860 $\pm$ 0.028 0.856 $\pm$ 0.029
-JointTrain (Upper) 0.913 $\pm$ 0.055 0.921 $\pm$ 0.045 0.908 $\pm$ 0.029 0.891 $\pm$ 0.041
+JointTrain 0.913 $\pm$ 0.055 0.921 $\pm$ 0.045 0.908 $\pm$ 0.029 0.891 $\pm$ 0.041
 Table 9: The Accuracy performance matrix of QPMIL-VL on reverse-order training.
@@
 To investigate the impact of different pooling operations on the performance of QPMIL-VL when generating query vector, we conduct comparative experiments using MaxPooling and MeanPooling.
-Table 10 presents the experimental results.
+Tab.
+10 presents the experimental results.
 The results show that QPMIL-VL performs better overall with MaxPooling compared to MeanPooling, particularly on forward-order training (the ACC increases from $0.877$$\pm$ $0.034$ to $0.890$$\pm$ $0.021$).